# Domain decomposition of the modified Born series approach for large-scale wave propagation simulations

### v2

@@ 3.1 Implementation
The domain decomposition framework is incorporated into the preconditioned Richardson iteration Eq. (6), but with the operators $L$ and $V$ computed for and applied to each subdomain. Thus, the fast convolution with ${(L+I)}^{-1}$ is computed for each subdomain $s$ as ${(L_{s}+I)}^{-1}$. It is the same for all subdomains of the same size and needs to be computed only once in that case. As discussed in Section 3, operator $V$ has three contributions, i.e. the scattering potential on the diagonal, which is now applied over the subdomains, the wrapping corrections $C$, and the transfer corrections (off-diagonal blocks of $A$). The steps for computing and applying the operator $V$ for every subdomain $s$ are as below:
@@
    Compute corrections ($C_{ss}\boldsymbol{x}_{s}$) for all edges based on the correction matrix $C_{ss}$ (described in Section 3). Note that $C_{ss}$ is a highly sparse matrix as the correction is truncated to $t\ll N$ points at the edge of the subdomain.
@@
    Transfer these edge corrections to the neighbouring subdomains.
@@
    Apply the scattering potential.
@@
    Add the edge corrections from the current subdomain, corresponding to the wrapping corrections $C_{ss}$.
@@
    Subtract the edge corrections coming from the neighbouring subdomains, which correspond to the transfer corrections $A_{s,s\pm 1}$, where the subscript $s\pm 1$ represents the neighbouring subdomains.
@@
Since we now use a different operator $V$ than the original MBS implementation, we can no longer use Eq. (4) to compute the scale factor $c$. Instead, we now require $ic<1/\left\lVert V\right\rVert$, where $\left\lVert V\right\rVert$ denotes the operator norm of $V$, i.e. its largest singular value [40]. For simplicity, we estimate this norm as the sum of the norms of all contributions to $V$: the original scattering potential, the wrapping correction, and domain transfer components, which gives an upper limit for $\left\lVert V\right\rVert$ and results in a scaling factor of
$c=-\frac{0.95i}{\left\lVert{k}^{2}(\boldsymbol{r})-k_{0}^{2}\right\rVert_{\infty}+\left(\sum_{d=1}^{3}a_{d}\right)\left\lVert A_{12}\right\rVert},$ (9)
where $a_{d}$ equals 00 if no wrapping or transfer correction is applied along dimension $d$ (corresponding to periodic boundaries). It equals $1$ if only wrapping corrections are needed (i.e. when there is only a single subdomain along this dimension) and $2$ if both the wrapping and transfer corrections are used. Thus, the wrapping and transfer corrections reduce the value of $c$, leading to slower convergence when $a_{d}\neq 0$. For a 1-domain simulation without wrapping corrections, $a_{d}=0$, and Eq. (9) reduces to Eq. (4).
Domain decomposition also offers an opportunity to improve the efficiency of the computations by initially activating only those subdomains that hold a source term. Neighbouring subdomains are activated only after the norm of the transferred field exceeds a certain threshold. Although this initially causes a small error, the MBS converges to the correct solution from any starting point and thus nullifies these errors.
With this activation strategy, for example, a 3D simulation of size $100\times 100\times 100$ wavelengths split into $3\times 1\times 1$ subdomains was $12\%$ faster than without this strategy.
@@ 4 Results and Discussion
In this section, we demonstrate the domain decomposition of the MBS in 3D simulations. For all simulations in this section, we simulate light propagation through a scattering structure illuminated with a planar source located at $x=0$. We use absorbing boundaries with a thickness of 5 wavelengths [3], and a sampling of 4 points per wavelength.
The structure consists of tightly packed spheres, with a relative wave velocity (or refractive index) of 1.33 and a small imaginary part, distributed randomly in a medium with a refractive index $n=1$. The spheres were generated using the random_spheres() function in PoreSpy [48]. Unless stated otherwise, we used spheres of radius $3$ wavelengths and a truncation size of $t=8$ in the simulations.
We start by validating the domain decomposition framework on a small sample and then analyse the accuracy and convergence of the domain decomposition approach for different truncation sizes $t$ and numbers of subdomains. Finally, we demonstrate domain decomposition in a large structure of $3.1\cdot 10^{7}$ cubic wavelengths.
@@ 4.1 Domain decomposition validation
[…]
Figure 5: Simulation results with and without domain decomposition for a 3D medium of size $50\times 50\times 50$ wavelengths. 2D slices at z/2 of (a) the refractive index distribution used in the simulation, (b) the result for a single domain, and (c) the result with three subdomains in the x direction, with the dash-dotted black lines indicating the subdomain boundaries. (d) Residual vs. iteration number for the simulations in b (solid blue line) and c (dashed black line).
Fig. 5a shows a slice, through the centre of the z-axis, of the 3D refractive index distribution used in this test. The results of the simulations are shown in Fig. 5b for the single domain and in Fig. 5c for the domain decomposed into three subdomains along the x-axis. The relative error between the two is just $1.7\cdot 10^{-4}$, confirming the validity of our domain decomposition approach.
Fig. 5d shows the residual as a function of the iteration number for the two simulations. It can be seen that the residual decreases monotonically, which is a key property of the MBS. The single-domain simulation converged in 409 iterations in 2.3 seconds, while the 3-domain simulation converged in 2,908 iterations in 11.7 seconds. The 7×7$\times$7 × increase in the number of iterations is due to a lower scaling coefficient $c$, with $a_{d}=2$ (Eq. (9)). Although this is a significant increase in the number of iterations, this factor is independent of the number of subdomains along an axis, as we will analyse in more detail in Section 4.2.2.
@@ 4.2.1 Dependence on truncation parameter
First, we examine the convergence behaviour as a function of $t$, the number of correction points used to generate the correction matrices $C_{11}$ and $A_{12}$ (Section 3). We decomposed the problem in Section 4.1 into two subdomains along the x-axis and varied $t$ from 00, indicating no correction, to $40$. We compare the relative error of these simulations with a single-domain simulation, as shown in Fig. 6a. The relative error between the two decays with more correction points, and with just $t=4$, the error is already less than $0.1\%$.
Figure 6: Convergence of a 3D simulation with two subdomains for different numbers of correction points $t$. (a) Relative error with a 1-domain simulation decays with $t$. (b) The number of iterations is almost constant for $t=4$ and higher, showing no dependence on $t$. (c) The total simulation time in seconds consistently increases for more correction points.
Fig. 6b shows the number of iterations for convergence as a function of the number of correction points, plotted for $t=4$ and higher, as the accuracy is low for $t<4$. The figure shows that the number of iterations does not depend on $t$. The time for convergence (Fig. 6c) consistently increases with more correction points, which is expected since the size of the correction matrix increases.
@@ 4.2.2 Dependence on number of subdomains
Next, we examine the number of iterations and time as a function of the number of subdomains (Fig. 7). Fig 7a shows the number of iterations for different numbers of subdomains along the x and y-axis. The lowest number of iterations (1160) is for a 1-domain simulation (black pixel in the lower left corner in Fig. 7a). This case corresponds to a scaling coefficient $c$ based on only a wrapping correction along the x-axis, for which the system is non-periodic. When there is more than one subdomain along either the x or y-axis, the scaling factor $c$ reduces, which increases the iterations to around $1751$ with more domains along the x-axis (dark pink band of pixels for $y=1$ in Fig. 7a), and around $2018$ for more domains along the y-axis (orange band of pixels for $x=1$ in Fig. 7a). For more than one subdomain in both the x and y direction, scaling coefficient $c$ is even lower, increasing the iterations to around $2448$ (Fig. 7a, the light yellow pixels for $x$ and $y>1$). Thus, the number of iterations increases only when subdomains are added along a new axis. The number of iterations is not dependent on the number of subdomains beyond one along an axis.
Figure 7: Number of iterations (a) and time (b) as a function of the number of subdomains along the x and y directions for a 3D simulation. The iteration count increases considerably only when subdomains are added along a new axis. The time increases with the number of subdomains but is consistently low for one domain in either x or y direction.
Fig. 7b shows the simulation time in seconds as a function of the number of subdomains along the x and y-axis. The simulation time is consistently low for one domain in x or y directions (black band of pixels along the x and y axes at $x=1$ or $y=1$). It consistently increases with more subdomains due to an increase in the amount of communication between them. The time is highest for $10$ domains along both x and y directions, i.e., the total number of subdomains = $100$ (top right corner of Fig. 7b).
@@ 4.3 Large scale simulations
To demonstrate the larger simulation size enabled by the domain decomposition approach, we run a simulation of $3.1\cdot 10^{7}$ cubic wavelengths on a system with an AMD EPYC 9354P CPU @ 3.25 GHz with 256 GB RAM and two GPUs (both NVIDIA L40S 48GB). The structure is the same as mentioned at the beginning of Section 4, except that the spheres have a radius of $6$ wavelengths and the clearance between them is set to $6$ wavelengths, which helps the energy propagate deeper into the simulation away from the source. The simulation size is a much larger $315\times 315\times 315$ wavelengths, corresponding to $1260\times 1260\times 1260$ voxels. Due to this large size, the simulation cannot be performed on a single GPU. Therefore, we run the single-domain simulation on the CPU and compare it against the 2-domain, 2-GPU simulation performed with the same system.
@@
Fig. 8a shows a slice, through the centre of the z-axis, of the 3D refractive index map used. The single-domain and 2-domain simulation results are shown in Figs. 8b and c, respectively. The insets in Fig. 8a-c show finer details in a zoomed-in region of the slices. The relative error between the two is $2.5\cdot 10^{-4}$.
@@
Figure 8: Large 3D simulation of size $315\times 315\times 315$ wavelengths with and without domain decomposition. A 2D slice of (a) the refractive index map, (b) the result from a 1-domain simulation on the CPU, and (c) the result from a 2-domain simulation using two GPUs, with the dash-dotted black line indicating the subdomain boundary. The insets are zoomed-in portions of the selected area to show finer details. (d) Residual as a function of the iterations for (b), a solid line, and (c), a dashed line.
@@
Fig. 8d shows the residual as a function of the iterations for the two simulations. The two-domain simulation converges in 9,878 iterations in 1.4 hours (4,958 s), while the single-domain simulation converges in 1,181 iterations in 18.2 hours (65,537 s) since it is run on the CPU (Table 1).
@@
In the 2-domain case, the wrapping and transfer corrections lower the scaling coefficient $c$, increasing the number of iterations by 8.3×8.3$\times$8.3 ×. In addition to this increase in the number of iterations, the domain decomposition framework has two overheads: the overhead of computing the edge corrections and the overhead due to data transfer between GPUs when the edge corrections are transferred and applied to neighbouring subdomains in each iteration. This subdomain communication leads to a lock-step execution, meaning the GPUs wait for the results of the other GPU(s). This overhead adds approximately $40\%$ time per cubic wavelength and reduces for larger simulations.
@@
Run parameters Total size (cubic wavelengths) Time (s) Iterations
2 domains, 2 GPUs $3.1\cdot 10^{7}$ $4,958.65$ $9,878$
1 domain, CPU $3.1\cdot 10^{7}$ $65,537.16$ $1,181$
Table 1: The largest possible simulation with domain decomposition, using 2 GPUs, and without, using the CPU. The simulations were run on a system with an AMD EPYC 9354P CPU @ 3.25 GHz with 256 GB RAM and two GPUs (both NVIDIA L40S 48GB). The domain decomposition simulation (row 1) takes 8.3×8.3$\times$8.3 × the number of iterations for convergence as the CPU simulation (row 2) but is 13.2×13.2$\times$13.2 × faster.
@@
Despite this overhead, the 2-domain, 2-GPU simulation is 13.2×13.2$\times$13.2 × faster than the single-domain simulation on the CPU without domain decomposition. With these two GPUs, we increased the simulation size by 1.93×1.93$\times$1.93 × of what can be accommodated on a single GPU to an unprecedented $315\times 315\times 315$ wavelengths. Furthermore, we showed that the number of domains can be increased arbitrarily. The overhead of adding additional domains is much smaller than the initial step from one to two subdomains, especially when the domain is split into subdomains along a single axis.

### v3

@@ 3.1 Implementation
The domain decomposition framework is incorporated into the preconditioned Richardson iteration (Eq. (6)), but with the operators $L$ and $V$ computed for and applied to each subdomain. Thus, the fast convolution with ${(L+I)}^{-1}$ is computed for each subdomain $s$ as ${(L_{ss}+I)}^{-1}$ (see Figs. 3b and 5b).
As discussed in Section 3, operator $V$ has three contributions, i.e., the scattering potential on the diagonal, the communication between the subdomains (off-diagonal blocks) and the wrapping artefacts with the opposite sign to $L$ in the diagonal blocks. The steps for applying the operator $V_{ss}$ for every subdomain $s$ are as below:
@@
    Compute the edge corrections (e.g., $A_{12}\boldsymbol{x}_{2}$) for all faces of all subdomains, using only $t$ voxels at the edges.
@@
    Transfer the edge corrections to neighbouring subdomains.
@@
    Apply the scattering potential, i.e., the diagonal elements of $V_{ss}$.
@@
    Subtract the edge corrections from the current subdomain (removes wrapping artefacts).
@@
    Add the edge corrections coming from the neighbouring subdomains (adds the field entering from the neighbouring subdomains).
The above procedure replaces the original implementation of $V$ and does not affect the iteration in Eq. (6) itself.
We use the relative residual $\epsilon\coloneqq\left\lVert\Gamma^{-1}\left(\boldsymbol{y}-A\boldsymbol{x}^{(n)}\right)\right\rVert_{2}/\left\lVert{\Gamma}^{-1}\boldsymbol{y}\right\rVert_{2}$ as a stopping criterion for the iteration. For the domain decomposition case, the global residual is given by $\sqrt{\sum_{s}\epsilon_{s}^{2}}$, where the summation is over all subdomains $s$. We use a threshold value of $10^{-6}$ as the stopping criterion. After the iteration has converged, the fields from the different subdomains are combined to give the final solution of the global field $\boldsymbol{x}$.
As we now use a different operator $V$ than the original MBS implementation, we can no longer use Eq. (4) to compute the scale factor $c$. We need to recompute $\left\lVert V\right\rVert$, the operator norm of $V$, for the domain decomposition case. For simplicity, we estimate this norm as the sum of the norms of all contributions to $V$: the original scattering potential, the subdomain communication blocks and the wrapping artefacts removal blocks, which gives an upper limit for $\left\lVert V\right\rVert$ and results in a scaling factor of
$c=-\frac{0.95i}{\left\lVert{k}^{2}-k_{0}^{2}\right\rVert_{\infty}+\left(\sum_{d=1}^{3}a_{d}\right)\left\lVert A_{12}\right\rVert},$ (9)
where $a_{d}$ equals 0 if no subdomain communication or removal of wrapping artefacts is applied along dimension $d$ (corresponding to periodic boundaries). It equals $1$ if only wrapping artefacts are removed (i.e., when there is only a single subdomain along this dimension and thus no subdomain communication) and $2$ if both the subdomain communication and removal of wrapping artefacts are applied. Thus, the subdomain communication and removal of wrapping artefacts reduces the value of $c$, leading to slower convergence when $a_{d}\neq 0$. It is important to note, however, that $a_{d}$ is independent of the number of domains along a dimension. So, increasing the number of subdomains along a dimension $d$ has no additional cost in terms of iterations. We discuss this in detail in Section 4.2.2. For a 1-domain simulation without removal of wrapping artefacts, $a_{d}=0$, and Eq. (9) reduces to Eq. (4).
Domain decomposition also offers an opportunity to improve the efficiency of the computations by initially activating only those subdomains that hold a source term. Neighbouring subdomains are activated only after the norm of the transferred field exceeds a certain threshold. Although this initially causes a small error, the MBS converges to the correct solution from any starting point and thus nullifies these errors. The number of iterations does not change due to this activation strategy, but the simulation time reduces, with the gains increasing with the number of domains and the simulation size. For example, a 3D simulation of size $100\times 100\times 100$ wavelengths split into $3\times 1\times 1$ subdomains was $12\%$ faster than without this strategy.
3.2 Complexity analysis
Table 1 compares the parameters of our domain decomposition method with the original MBS [2], a recent FFT-accelerated solver for the wave equation [54], PSTD [10], and FDTD methods [6, 7].
The accuracy of our domain decomposition approach, the original MBS, and the FFT-accelerated solver by Vico et al. [54] is only restricted by machine precision, whereas that of FDTD is proportional to $\Delta x^{2}$ and $\Delta t^{2}$, and that of PSTD is proportional to $\Delta t^{2}$ due to the finite different approximations. Although our domain decomposition method and the MBS have no time dependence, we can interpret the ‘pseudo-propagation’ length [2], i.e. the distance the solution expands with each iteration, to the real propagation steps of the PSTD and FDTD methods. The pseudo-propagation distance, given by $\lvert 2k_{0}c\rvert$, depends on the scaling parameter $c$ for both the original MBS (see Eq. (4)) and our domain decomposition method (see Eq. (9)).
The computational complexity of the original MBS implementation has been studied in literature [2, 4], and it is dominated by the computation of the FFT in the MBS algorithm. For a $D$-dimensional problem with $N^{D}$ voxels, the method requires $\mathcal{O}\left(N^{D}\textrm{log}N^{D}\right)$ floating point operations per iteration, the same as the FFT-accelerated solver by Vico et al. [54] and the PSTD method [10].
For the domain decomposition method, we divide the $N^{D}$ voxels over $M$ total subdomains. Then we need to compute a total of $DM$ FFTs, each over a subdomain of size $N^{D}/M$, giving a complexity of $\mathcal{O}\left(N^{D}\textrm{log}(N^{D}/M)\right)$ for the FFTs. The overhead due to the domain decomposition is $\mathcal{O}\left(MDN^{D-1}t\right)$, which is asymptotically smaller than that of the FFT.
This work MBS [2] Vico et al. [54] PSTD [10] FDTD [6, 7]
Grid spacing $\Delta x$ $\leq\lambda_{m}/2$ $\leq\lambda_{m}/8$
Accuracy tolerance adjustable $\mathcal{O}(\Delta t^{2})$ $\mathcal{O}(\Delta t^{2}\Delta x^{2})$
to machine precision
Time step $\Delta t$ n/a $\leq\frac{2\Delta x}{\pi v_{c}\sqrt{D}}$ $\leq\frac{\Delta x}{v_{c}\sqrt{D}}$
(Pseudo-) propagation $\lvert 2k_{0}c\rvert$ no estimate available $\frac{2}{k_{m}\sqrt{D}}$ $\frac{\pi}{4k_{m}\sqrt{D}}$
FLOPs per iteration $\mathcal{O}\left(N^{D}\textrm{log}\frac{N^{D}}{M}\right)$ $\mathcal{O}(N^{D}\textrm{log}N^{D})$ $\mathcal{O}(N^{D})$
@@ 3.2 Complexity analysis
Table 1: Key parameters for our domain decomposition approach (“This work”), compared to the original MBS [2], the fast Fourier transform (FFT)-accelerated solver in [54], the PSTD [10] and FDTD [6, 7] methods, following Table 1 from Osnabrugge et al. [2]. Here, $k_{m}\coloneqq\left\lVert k^{2}\right\rVert_{\infty}$ is the largest wavenumber, $\lambda_{m}\coloneqq 2\pi/k_{m}$ is the shortest wavelength, $D$ is the dimensionality of the system, $v_{c}$ is the speed of light in the simulated medium, and $T$ is the time needed for a wave to propagate through the full simulation.
3.3 Memory usage
The memory usage of the MBS method is approximately $40$ bytes per voxel. The factor $40$ originates from $5$ arrays made up of complex floats of $8$ bytes each, namely, the refractive index, the ${(L+I)}^{-1}$ operator, the (IV)(I-V) operator (Eq. (6)), an array for updating and storing the field $\boldsymbol{x}$, and an internal FFT array.
On top of this, our domain decomposition strategy requires approximately $8\cdot 2\cdot t$ bytes per voxel at each subdomain interface, with $8$ bytes per complex float and $t$ the truncation parameter. We use a default value of $t=8$ in our simulations, so the relation reduces to $8\cdot 2\cdot 8=128$ bytes per voxel at each subdomain interface.
For instance, we can compute the memory usage for the $320\times 320\times 320$ wavelengths structure sampled at $4$ points per wavelength that we later present results for in Section 4.3. We use absorbing boundaries of $5$ wavelengths along the x-axis and periodic boundaries along the y and z axes, giving the resultant simulation size of $330\times 320\times 320$ wavelengths, or $1320\times 1280\times 1280$ voxels ($2.16\times 10^{9}$ voxels). The memory needed for a one-domain simulation is $40\cdot 1320\cdot 1280^{2}=86.5$ GB. The largest VRAM with a single GPU available to us is $48$ GB with the NVIDIA A40 GPUs, so we split the problem over two GPUs into two domains along the x-axis, and there are $1280^{2}$ voxels at the subdomain interface. The overhead is calculated as only $128\cdot 1280^{2}=0.2$ GB, giving a total memory use of $86.7$ GB for the largest simulation we could run on two GPUs with a total memory of $96$ GB, indicating an unaccounted overhead of around $9.7\%$ in the memory usage calculations.
@@ 4 Results and Discussion
In this section, we demonstrate the domain decomposition of the MBS in 3D simulations. For all simulations in this section, we simulate light propagation through a scattering structure illuminated with a planar source located at $x=0$. We use absorbing boundaries [3] with a thickness of 5 wavelengths along the x-axis and periodic boundaries along the y and z axes. We use a sampling of 4 points per wavelength, and a threshold value of $10^{-6}$ for the relative residual as the stopping criterion (see Section 3.1). In all the simulations, we split the problem into subdomains of equal size.
The structure consists of tightly packed spheres, with a refractive index of $1.33$ and a small imaginary part of $0.01$, distributed randomly in a medium with a refractive index of $1$. The spheres were generated using the random_spheres() function in PoreSpy [55]. Unless stated otherwise, the spheres have a radius of $3$ wavelengths and the truncation parameter $t=8$ in the simulations.
We first confirmed that our domain decomposition method converges to the analytical solution for propagation through an empty medium. In this section, we start by validating the domain decomposition framework on a small sample and then analyse the accuracy and convergence of the domain decomposition approach for different values of the truncation parameter $t$ and numbers of subdomains. Finally, we demonstrate domain decomposition in a large structure of $3.27\cdot 10^{7}$ cubic wavelengths, a size that takes considerably longer to run on a CPU but can be completed significantly faster by splitting the problem over two GPUs.
@@ 4.1 Domain decomposition validation
We first validate the accuracy of our domain decomposition framework by comparing it to the single-domain MBS approach for a 3D problem of size $50\times 50\times 50$ wavelengths. With absorbing boundaries [3] of $5$ wavelengths along the x-axis, the total simulation size is $60\times 50\times 50$ wavelengths. We quantify the accuracy by comparing the field obtained from domain decomposition ($\boldsymbol{x}$) and the single-domain simulation result ($\boldsymbol{x}_{\mathrm{ref}}$), excluding the absorbing boundaries, with the relative error $\left\lVert\boldsymbol{x}-\boldsymbol{x}_{\mathrm{ref}}\right\rVert_{2}^{2}/\left\lVert\boldsymbol{x}_{\mathrm{ref}}\right\rVert_{2}^{2}$.
Figure 6: Simulation results with and without domain decomposition for a 3D medium of size $50\times 50\times 50$ wavelengths. 2D slices at z/2 of (a) the refractive index distribution used in the simulation, (b) the result for a single domain, and (c) the result with three subdomains in the x direction, with the dash-dotted black lines indicating the subdomain boundaries. The spheres appear to be of different sizes because the 2D slice takes a cross-section of the spheres at different points. (d) Residual vs. iteration number for the simulations in b (solid blue line) and c (dashed black line).
Fig. 6a shows a slice through the centre of the z-axis of the 3D refractive index distribution used in this test, without absorbing boundaries. The results of the simulations are shown without absorbing boundaries in Fig. 6b for the single domain of simulation size $60\times 50\times 50$ wavelengths and in Fig. 6c for the domain decomposed into three subdomains along the x-axis, each of size $20\times 50\times 50$ wavelengths. The relative error between the two is just $2\cdot 10^{-4}$, confirming the validity of our domain decomposition approach.
Fig. 6d shows the residual as a function of the iteration number for the two simulations. It can be seen that the residual decreases monotonically, which is a key property of the MBS and is also conserved in our domain decomposition method. The single-domain simulation converges (i.e., reaches a relative residual of $10^{-6}$) in 584 iterations in 6.7 seconds, while the 3-domain simulation converges in 1751 iterations in 8.1 seconds. The $3\times$ increase in the number of iterations is due to a lower scaling coefficient $c$, with $a_{d}=2$ (Eq. (9)). Although this is a significant increase in the number of iterations, this factor is independent of the number of subdomains along an axis, as we analyse in more detail in Section 4.2.2.
@@ 4.2.1 Dependence on truncation parameter
First, we examine the convergence behaviour as a function of the truncation parameter $t$ used to generate the blocks for subdomain communication and removal of wrapping artefacts. We decompose the problem in Section 4.1 into two subdomains of equal size along the x-axis ($30\times 50\times 50$ wavelengths each) and vary $t$ from 0 to $40$. We compare the relative error of these simulations with a single-domain simulation, as shown in Fig. 7a. The relative error between the two decays with increasing $t$, and with just $t=4$, the error is already less than $0.1\%$.
Figure 7: Convergence of a 3D simulation with two subdomains for different values of the truncation parameter $t$. (a) Relative error with a 1-domain simulation decays with $t$. (b) The number of iterations is almost constant for $t=4$ and higher, showing no dependence on $t$. (c) The total simulation time in seconds consistently increases with $t$.
Fig. 7b shows the number of iterations for convergence as a function of $t$, plotted for $t=4$ and higher, as the accuracy is low for $t<4$. The figure shows that the number of iterations does not depend on $t$. The time for convergence (Fig. 7c) consistently increases with $t$, which is expected as the size of the blocks for subdomain communication and removal of wrapping artefacts increases.
@@ 4.2.2 Dependence on number of subdomains
Next, we examine the number of iterations and time as a function of the number of subdomains (Fig. 8). Fig 8a shows the number of iterations for different numbers of subdomains along the x and y-axis. The lowest number of iterations (1160) is for a 1-domain simulation (black pixel in the lower left corner in Fig. 8a). This case corresponds to a scaling coefficient $c$ based on only the removal of wrapping artefacts along the x-axis, for which the system is non-periodic. When there is more than one subdomain along either the x or y-axis, the scaling factor $c$ reduces, which increases the iterations to around $1751$ with more domains along the x-axis (dark pink band of pixels for $y=1$ in Fig. 8a), and around $2018$ for more domains along the y-axis (orange band of pixels for $x=1$ in Fig. 8a). For more than one subdomain in both the x and y direction, the scaling coefficient $c$ is even lower, increasing the iterations to around $2448$ (Fig. 8a, the light yellow pixels for $x$ and $y>1$). Thus, the number of iterations increases only when subdomains are added along a new axis. The number of iterations is not dependent on the number of subdomains beyond one along an axis, indicating the scalability of our domain decomposition framework. This result also follows from the modified scaling factor $c$ in Eq. (9), which has no dependence on the number of subdomains.
Figure 8: Number of iterations (a) and time (b) as a function of the number of subdomains along the x and y directions for a 3D simulation. The iteration count increases only when subdomains are added along a new axis, which follows from the scaling factor $c$ in Eq. (9) being independent of the number of subdomains. The time increases with the number of subdomains, but is consistently low for one domain in either x or y direction. Note that in both (a) and (b), the colour axis does not start from 0.
Fig. 8b shows the simulation time in seconds as a function of the number of subdomains along the x and y-axis. The simulation time is consistently low for one domain in the x or y directions (black band of pixels along the x and y axes at $x=1$ or $y=1$). It consistently increases with more subdomains due to an increase in the amount of communication between them. The time is highest for $10$ domains along both x and y directions, i.e., the total number of subdomains = $100$ (top right corner of Fig. 8b).
@@ 4.3 Large scale simulations
To demonstrate the larger simulation size enabled by the domain decomposition approach, we solve a problem of $3.27\cdot 10^{7}$ cubic wavelengths ($320\times 320\times 320$ wavelengths) on a system with two Silver-4216 2.10 GHz CPUs with 128 GB RAM and four A40 48GB GPUs. The structure is the same as mentioned at the beginning of Section 4, except that the spheres have a radius of 6 wavelengths and the clearance between them is set to 6 wavelengths, which helps the energy propagate deeper into the simulation away from the source. The simulation size is $330\times 320\times 320$ wavelengths (with absorbing boundaries of $5$ wavelengths along the x-axis), corresponding to $1320\times 1280\times 1280$ voxels ($2.16$ Gigavoxels). Due to this large size, the simulation cannot be performed on a single GPU. It can, however, be run on a CPU, with a significant tradeoff in terms of computation time. Therefore, we run the single-domain simulation on the CPU and compare it against the 2-domain, 2-GPU simulation performed with the same system. Each of the two subdomains is of size $165\times 320\times 320$ wavelengths. We also run the same simulation on 3 and 4 GPUs, split over as many subdomains (over the x dimension), to demonstrate the scalability of the domain decomposition framework.
@@
Fig. 9a shows a slice of the 3D refractive index map through the centre of the z-axis without absorbing boundaries. The single-domain and 2-domain simulation results are shown in Figs. 9b and c, respectively. The relative error between the two is $2.9\cdot 10^{-4}$. The insets in Fig. 9a-c show finer details in a zoomed-in region of the slices.
@@
Figure 9: Large 3D simulation of size $320\times 320\times 320$ wavelengths with and without domain decomposition. A 2D slice of (a) the refractive index map, (b) the result from a 1-domain simulation on the CPU, and (c) the result from a 2-domain simulation using two GPUs, with the dash-dotted black line indicating the subdomain boundary. The insets are zoomed-in portions of the selected area to show finer details. The spheres appear to be of different sizes because the 2D slice takes a cross-section of the spheres at different points. (d) Residual as a function of the iterations for (b), a solid line, and (c), a dashed line.
@@
Fig. 9d shows the residual as a function of the iterations for the two simulations. The two-domain simulation (Fig. 9c) that is run on 2 GPUs converges in 4697 iterations in $45$ minutes (2730 s), while the single-domain simulation (Fig. 9b) converges in 1316 iterations in $15.5$ hours (56040 s) as it is run on the CPU (Table 2).
@@
In the 2-domain case, the subdomain communication and removal of wrapping artefacts lower the scaling coefficient $c$, increasing the number of iterations by $3.5\times$. The 2-domain simulation is $20\times$ faster than the 1-domain simulation on the CPU. However, the overhead of the iteration increase is a trade-off to be considered. In addition to this increase in the number of iterations, the domain decomposition framework has two overheads: the overhead of computing the edge corrections and the overhead due to data transfer between GPUs when the edge corrections are transferred and applied to neighbouring subdomains in each iteration. This subdomain communication leads to a lock-step execution, meaning the GPUs wait for the results of the other GPU(s). This overhead adds approximately $40\%$ time per cubic wavelength and reduces for larger simulations.
@@
Run parameters Total size (cubic wavelengths) Time (s) Time $\cdot$ #GPU (s) Iterations
1 domain, CPU $3.27\cdot 10^{7}$ 56040 1316
2 domains, 2 GPUs $3.27\cdot 10^{7}$ 2730 5460 4697
3 domains, 3 GPUs $3.27\cdot 10^{7}$ 2022 6066 4697
4 domains, 4 GPUs $3.27\cdot 10^{7}$ 1600 6400 4697
Table 2: The largest possible simulation with domain decomposition on 2 GPUs, with simulation runs on the CPU for comparison, and 3 and 4 GPUs to illustrate scalability. The simulations were run on a system with two Silver-4216 2.10 GHz CPUs with 128 GB RAM and four A40 48GB GPUs. The domain decomposition simulations on the GPUs (rows 2-4) take $3.5\times$ the number of iterations for convergence as the CPU simulation (row 1), but are at least $20\times$ faster. Domain decomposition reduces the total computation time (column 3), indicating parallel execution, and adds only about 10% overhead to the total amount of GPU-seconds (column 4) for each GPU added, while leaving the total number of iterations constant (column 5) from 2 domains onwards.
@@
Despite this overhead, the 2-domain, 2-GPU simulation is $20\times$ faster than the single-domain simulation on the CPU without domain decomposition. With these two GPUs, we increased the simulation size by $1.95\times$ of what can be accommodated on a single GPU to an unprecedented $320\times 320\times 320$ wavelengths. When the domain is split into more subdomains along a single axis, there is no additional overhead in terms of iterations beyond the initial step of going from one to two subdomains, as seen from the last column in Table 2. When more GPUs are used, the total simulation time reduces, indicating a parallelisation of the computations. The total number of GPU-seconds used for the simulation only increases by about 10% overhead per GPU, indicating a small synchronisation and communication overhead.